# Data Integration with Fusion Searchlight: Classifying Brain States from Resting-state fMRI

### v1

@@ Abstract
Spontaneous neural activity observed in resting-state fMRI is characterized by complex spatio-temporal dynamics. Different measures related to local and global brain connectivity and fluctuations in low-frequency amplitudes can quantify individual aspects of these neural dynamics. Even though such measures are derived from the same functional signals, they are often evaluated separately, neglecting their interrelations and potentially reducing the analysis sensitivity. In our study, we present a fusion searchlight (FuSL) framework to combine the complementary information contained in different resting-state fMRI metrics and demonstrate how this can improve the decoding of brain states. Moreover, we show how explainable AI allows us to reconstruct the differential impact of each metric on the decoding, which additionally increases spatial specificity of searchlight analysis. In general, this framework can be adapted to combine information derived from different imaging modalities or experimental conditions, offering a versatile and interpretable tool for data fusion in neuroimaging.
@@ 1 Introduction
Resting-state fMRI (rs-fMRI) studies intrinsic neural activity of the brain at rest and is characterized by complex spatial and temporal dynamics (Biswal et al., 1995). Until now, numerous methods have been established to describe and analyze different aspects of these spontaneous neural dynamics. Functional connectivity (FC) is a popular approach for studying global interactions between brain regions from a functional network perspective (van den Heuvel and Hulshoff Pol, 2010). Measures from graph theory allow us thereby to detect changes in connectivity properties of brain networks (Bassett and Sporns, 2017). On a smaller scale, regional homogeneity (ReHo) quantifies local connectivity across the cortex in the range of millimeters, and is defined as the temporal coherence or synchrony of the BOLD signal of neighboring voxels or vertices (Jiang and Zuo, 2016). In addition to connectivity measures, the fractional amplitude of low frequency fluctuations (fALFF) can characterize local modulations in the amplitude of signals (Yang et al., 2007, Zou et al., 2008). This quantity reflects the ratio of the BOLD signal’s power spectrum of low frequencies to the signal’s entire frequency range. In this manner, these different rs-fMRI metrics capture complementary aspects of BOLD signal dynamics (Lv et al., 2018). In clinical MRI studies for instance, changes in neural dynamics can be reflected simultaneously in multiple metrics. In rs-fMRI studies on depression, alterations have been linked to modulations in FC (Javaheripour et al., 2021, Kaiser et al., 2015), fALFF (Gao et al., 2021, Wang et al., 2022) and ReHo (Shen et al., 2017, Ni et al., 2023, Sun et al., 2022). Also multiple sclerosis has shown to affect a multitude of markers, such as FC (Tahedl et al., 2018, Huang et al., 2020, Tahedl et al., 2022), fALFF (Liu et al., 2015) and ReHo (Wu et al., 2016). Furthermore, in a pharmacological rs-fMRI study, the sedative effects of alprazolam were observed simultaneously in the modulations of FC, ReHo and fALLF (Wein et al., 2024). However, many current rs-fMRI studies focus only on a single metric, or analyze these biomarkers separately. In this manner, information on important modulations of the BOLD signal may be neglected, potentially reducing sensitivity of detecting altered brain states [^1] Based on the definition of Greene et al. (2023), we refer as brain states to recurring activity patterns distributed across the brain that emerge from physiological or cognitive processes..
@@
In our study we present an interpretable framework based on searchlight (SL) decoding for fusing such complementary fMRI markers and improving analysis sensitivity. SL decoding uses multivariate pattern analysis (MVPA) to test the information content in small spherical subsets in every voxel or vertex across the cortex (Kriegeskorte et al., 2006, Etzel et al., 2013). In this manner SL typically produces spatial maps of decoding accuracies, highlighting areas that are informative for a specific brain state. In our fusion SL (FuSL) framework we integrate the distinct information of BOLD dynamics expressed in local ReHo, fALFF and global FC and show how this increases decoding performance in rs-fMRI. In a next step we utilize an explainable AI (XAI) approach to retrospectively trace back the feature importance of each metric for decoding. To this aim, we rely on Shapley additive explanations (SHAP) (Lundberg and Lee, 2017), a model-agnostic XAI method based on game theoretically optimal Shapley values (Shapley, 1953). Moreover, we show that incorporating XAI to evaluate feature importance has the additional benefit of increasing spatial specificity in SL analysis. Solely analyzing prediction accuracy, as typically done in current studies applying SL decoding, only reveals if there is somewhere within the SL radius information present. Thus, maps of decoding accuracies do not exactly represent maps of informative voxels (Etzel et al., 2013). Incorporating SHAP as a post hoc test can help us to reconstruct the precise locations of voxels within a SL that are most important for decoding, what helps us to increase spatial specificity of SL analysis.
@@
We first investigate general properties the FuSL framework employing an artificial MRI dataset and study how different combinations of informative and uninformative data sources affect its decoding performance. With rs-fMRI data from a pharmacological fMRI study we further demonstrate that the integration of complementary rs-fMRI metrics can enhance the prediction of pharmacological treatment of participants. Sedative effects of benzodiazepines have shown to be reliably expressed in different rs-fMRI markers, such as FC (Wei et al., 2013, Hashmi et al., 2017, Schrouff et al., 2011), ReHo (Blanco-Hinojo et al., 2021) and fALFF (Kiviniemi et al., 2003, 2005). We show how combining these metrics can help us to detect sedation-related alterations in the BOLD signal, based on rs-fMRI data from participants who received either a benzodiazepine (alprazolam) or a placebo during a pharmacological study (Wein et al., 2024). We finally demonstrate how SHAP can be used to trace back the differential impact of sedation on the individual fMRI markers.
Numerous data fusion frameworks have meanwhile shown to be beneficial in diverse areas of neuroimaging analyses. In clinical neuroscience integrating distinct modalities such as functional MRI, structural MRI, EEG and MEG has improved our understanding on mental disorders (Calhoun and Sui, 2016). Fusing MEG or EEG with fMRI data using representational similarity analysis (RSA) allows characterizing neural activation simultaneously with high temporal and spatial resolution in cognitive neuroscience (Cichy et al., 2016, Cichy and Oliva, 2020). In brain connectivity research adding structural information to functional MRI data has helped adding biological constraints to computational and machine learning models of neural dynamics (Cabral et al., 2014, Suárez et al., 2020, Wein et al., 2021b, a). Merging information expressed in different structural MRI modalities has enhanced accuracy of segmenting tumors (Ahamed et al., 2023) and multiple sclerosis lesions (Zeng et al., 2020). In this study we demonstrate that data fusion can be also beneficial for SL decoding analysis. Besides improving decoding in rs-fMRI, our FuSL framework can be easily adapted to integrate information from different imaging modalities. While deep learning frameworks often require a large amount of training data and manual fine-tuning of model parameters (Schulz et al., 2020), SL has proven to be a robust and widely-used MVPA tool that limits its complexity inherently by using spatially localized information. Furthermore, using a model-agnostic feature attribution method based on Shapley values allows one to use a wide range of MVPA models for decoding in our workflow. In this manner our FuSL framework can provide a flexible and interpretable framework for diverse data fusion applications in neuroimaging.
2 Methods and Materials
2.1 Fusion searchlight
In our FuSL framework we integrate the complementary information provided by different data sources in a searchlight-based decoding analysis. For decoding a brain state based on resting-state fMRI we first computed regional homogeneity (ReHo), fractional amplitude of low frequency fluctuations (fALFF) and regional functional connectivity efficiency (FCE), and concatenated these values within each SL to train a classifier (figure 1 A and B). After training and testing the SL classifier using stratified cross validation (CV) (Kohavi, 1995), we obtained a spatial map of test accuracies across the cortex, pointing out vertices with neighborhoods that were informative to decode a specific brain state, like in classical SL analysis (figure 1 (C)). To identify test accuracies that were significantly different from chance level we permuted the labels and applied threshold-free cluster enhancement (TFCE) to the original an permuted accuracy maps (Smith and Nichols, 2009). To correct for multiple comparisons across vertices we took the maximum TFCE values of accuracy maps derived from permuted data to build the null distribution and tested the values of the original TFCE map against them (Smith and Nichols, 2009).
In a next step we aimed to explain the predictions of the FuSL classifier. To reveal which metrics had driven the decision of the SL, we used Shapley additive explanations (SHAP) to reconstruct the importance of each metric at each location on the cortex (Lundberg and Lee, 2017). This explanation approach assigns post-hoc an importance score to each input feature of a classifier, which we utilized to determine whether modulations in ReHo, fALFF or FCE were indicative for decoding a specific brain state. For each SL we first computed the feature importance of each metric at each vertex and then averaged these values across overlapping SLs at each vertex. In this manner we obtained one importance score for each metric at each vertex on the cortex. A formal derivation of these importance measure is provided in Supplement I. To obtain maps that reflect simultaneously the importance and the informativeness of a metric, we further weighted spatial maps of SHAP values (feature importance) with the spatial map of decoding accuracies (informativeness). We denoted this combination of informativeness and importance as the actual impact of a metric (figure 1 (D)). However, these maps do not provide us with information whether a metric is increased or decreased due to an intervention or condition, which can be valuable information for the interpretation of analysis results. To address this shortcoming we propose to additionally incorporate a third characteristic of the data. We computed the differences in metrics between the groups across the cortex to obtain a spatial map indicating the intervention-associated changes of each metric. Finally we weighted these maps of feature differences with the feature impact. This yielded spatial maps that condense the modulation, importance and informativeness of each metric at each location for a compact visual interpretation. We denote this combination as the feature-weighted impact. For a more specific analysis of these three aspects one can also visualize these maps individually, depending on the requirements of MRI study. An implementation including a demo version is available at: https://github.com/simonvino/FuSL which is based on functionalities provided by NiLearn [^2] https://nilearn.github.io/stable/index.html (Abraham et al., 2014), scikit-learn [^3] https://scikit-learn.org/stable/ (Pedregosa et al., 2011), and the SHAP software package [^4] https://shap.readthedocs.io/en/latest/ (Lundberg and Lee, 2017).
@@ 2.2.1 Artificial dataset
In the beginning we use an artificial dataset to investigate basic properties of FuSL when combining different data sources. On one side we want to study decoding performance of FuSL when signal or information is present in multiple imaging metrics or modalities simultaneously in the same location. On the other side we explore the behavior when we combine noisy non-informative data sources with informative ones in our FuSL. To study these scenarios we generate 3 artificial source signals, which could represent 3 different neuroimaging metrics or modalities. In this toy example the objective of the FuSL is, based on these 3 sources, to discriminate between two groups. We generated data for two groups, each containing 30 samples, by inducing artificial signals in one group, but not the other. Therefore we set signal values to 1 within a circular region of interest (ROI). In source 1 and 2 we simultaneously induce a signal in the left frontal cortex setting values within a 5 mm radius to 1 and -1 respectively (defined as ROI 1). As a third source we induced a signal in the left parietal cortex, setting values to 1 within a 5 mm radius (definded as ROI 2). We then generated multiple samples from these signals by sampling the signal amplitudes from a Gaussian normal distribution with standard deviation of 1, to simulate signal variability as observed in biological fMRI data. We additionally added random Gaussian noise to each sample with varying powers $P_{noise}=[0.1,1,2]$ resulting in three different datasets with signal to noise ratios $S/N=[2,1,0.1]$ respectively. In this manner we created datasets for which signals in source 1 and 2 overlapped in ROI 1, and one signal was only expressed in ROI 2 in source 3.
@@ 2.2.2 Resting-state fMRI dataset
We used rs-fMRI data of 34 subjects who underwent rs-fMRI after a five days intake of either alprazolam or placebo (Wein et al., 2024). Participants gave their written informed consent at the beginning of the experiment and the study was approved by the local ethics committee (approval number 18-1197-111) and the National Institute for Pharmaceutical Security (BfArM, Eudra-CT-number: 2018-002181-40). Subjects received alprazolam (1.5 mg/d in 3 doses of 0.5 mg) or placebo for 5 days, each with at least 7 days washout phases between treatments.
@@
For preprocessing the functional and structural imaging data, we incorporated the processing pipeline provided by fMRIPrep [^5] https://fmriprep.org/en/stable/ (version 20.2.4) (Esteban et al., 2017) which included bias field correction, motion correction, slice timing correction and projection onto the cortical surface, generating high-resolution grayordinate time courses in the fsLR standard space (Glasser et al., 2013). Prior to the ReHo and fALFF analysis we smoothed the fMRI data using Gaussian surface smoothing with a FWHM of 3 mm (Glasser et al., 2013).
@@
For computing regional homogeneity (ReHo), we selected for each vertex its 5-hop neighboring vertices in the fsLR standard space (Glasser et al., 2013), and computed ReHo as the average Pearson correlation between neighbouring BOLD time courses within the 0.01 Hz - 0.1 Hz frequency range. To analyze fractional amplitude of low frequency fluctuations (fALFF) we focused on the very-low frequency range 0.01 Hz - 0.05 Hz, which has been reported to increase during midazolam sedation and anesthesia (Kiviniemi et al., 2000, 2003, 2005). We computed fALFF as the ratio of the BOLD signal’s power spectrum of these very low-frequencies to the signal’s entire frequency range. To evaluate whole-brain FC we subdivided each hemisphere into $180$ ROIs defined by Glasser et al. (2016). We computed the average activity time course within each ROI and filtered the resting-state BOLD signal within the 0.01 Hz - 0.1 Hz low frequency range. FC was then defined as the Pearson correlation coefficient $r$ between time courses of pairs of ROIs to obtain a $360\times 360$ whole-brain FC matrix. We selected a moderate FC threshold $r\geq 0.6$ yielding an average connection density of $39.2\%$ and analyzed modulations in functional connectivity efficiency (FCE) for each ROI in the network. According to the weights of the probabilistic multi-modal parcellation (Glasser et al., 2016), for each vertex we computed FCE as a weighted sum of the ROI-wise FCE values, weighting each vertex with the probability to belong to a respective ROI. This yielded a spatially continuous distribution of FCE values across vertices. A mathematical definition of these three metrics is additionally provided in supplement I.

### v2

@@ Abstract
Resting-state fMRI captures spontaneous neural activity characterized by complex spatiotemporal dynamics. Various metrics, such as local and global brain connectivity and low-frequency amplitude fluctuations, quantify distinct aspects of these dynamics. However, these measures are typically analyzed independently, overlooking their interrelations and potentially limiting analytical sensitivity. Here, we introduce the Fusion Searchlight (FuSL) framework, which integrates complementary information from multiple resting-state fMRI metrics. We demonstrate that combining these metrics enhances the accuracy of pharmacological treatment prediction from rs-fMRI data, enabling the identification of additional brain regions affected by sedation with alprazolam. Furthermore, we leverage explainable AI to delineate the differential contributions of each metric, which additionally improves spatial specificity of the searchlight analysis. Moreover, this framework can be adapted to combine information across imaging modalities or experimental conditions, providing a versatile and interpretable tool for data fusion in neuroimaging.
@@ 1 Introduction
Complex cognition arises from the interplay of activity among billions of neurons. A fundamental goal in studying neurological diseases and mental health disorders is to understand how these interactions may be systematically altered to explain behavioral changes. One promising noninvasive approach involves studying the brain at rest (Biswal et al., 1995) and characterizing brain activity using measures of local BOLD activity, as well as localized and distributed connectivity metrics. These metrics can be compared between clinical groups and healthy controls or between a baseline and a post-intervention state.
Resting-state fMRI markers
@@
Functional connectivity (FC) is one popular approach for studying global interactions between brain regions from a functional network perspective (van den Heuvel and Hulshoff Pol, 2010). Thereby, measures from graph theory allow us to detect changes in connectivity properties of brain networks (Bassett and Sporns, 2017). On a smaller scale, regional homogeneity (ReHo) quantifies local connectivity across the cortex in the range of millimeters, and is defined as the temporal coherence or synchrony of the BOLD signal of neighboring voxels or vertices (Jiang and Zuo, 2016). In addition to connectivity measures, the fractional amplitude of low frequency fluctuations (fALFF) can characterize local modulations in the amplitude of signals (Yang et al., 2007, Zou et al., 2008). In this manner, these different resting-state fMRI (rs-fMRI) metrics capture complementary aspects of BOLD signal dynamics (Lv et al., 2018).
In clinical rs-fMRI studies different markers can be modulated simultaneously
In rs-fMRI studies on depression, alterations have been linked to modulations in FC (Javaheripour et al., 2021, Kaiser et al., 2015), fALFF (Gao et al., 2021, Wang et al., 2022) and ReHo (Shen et al., 2017, Ni et al., 2023, Sun et al., 2022). Also multiple sclerosis has shown to affect a multitude of markers, such as FC (Tahedl et al., 2018, Huang et al., 2020, Tahedl et al., 2022), fALFF (Liu et al., 2015) and ReHo (Wu et al., 2016). Furthermore, sedative related effects have shown to be apparent across multiple different rs-fMRI metrics. Administering chloral hydrate, dexmedetomidine and propofol resulted in decreases in whole-brain FC density and efficiency (Wei et al., 2013, Hashmi et al., 2017, Schrouff et al., 2011). Intake of alprazolam increased ReHo in sensory- and sensory-integration areas (Blanco-Hinojo et al., 2021). Administering diazepam (Pflanz et al., 2015) and midazolam (Liang et al., 2015) yielded an increase of within-network connectivity mainly in low-level sensory networks. Furthermore, midazolam elevated low frequency amplitudes ($<0.05Hz$) in the visual, sensorimotor and auditory resting-state network (Kiviniemi et al., 2003, 2005). However, many current rs-fMRI studies focus only on a single metric, or analyze these biomarkers separately. In this manner, information on important modulations of the BOLD signal may be neglected, potentially reducing sensitivity of detecting altered brain states [^1] Based on the definition of Greene et al. (2023), we refer as brain states to recurring activity patterns distributed across the brain that emerge from physiological or cognitive processes..
Integrating rs-fMRI metrics with fusion searchlight
@@
In our study we present an interpretable framework based on searchlight (SL) decoding for fusing such complementary fMRI markers and improving analysis sensitivity. SL decoding uses multivariate pattern analysis (MVPA) to test the information content in small spherical subsets in every voxel or vertex across the cortex (Kriegeskorte et al., 2006, Etzel et al., 2013). In this manner SL typically produces spatial maps of decoding accuracies, highlighting areas that are informative for a specific brain state. In our fusion SL (FuSL) framework we integrate the distinct information of BOLD dynamics expressed in local ReHo, fALFF and global FC and show how this increases decoding performance in rs-fMRI. In a next step we utilize an explainable AI (XAI) approach to retrospectively trace back the feature importance of each metric for decoding. To this aim, we rely on Shapley additive explanations (SHAP) (Lundberg and Lee, 2017), a XAI method based on game theoretically optimal Shapley values (Shapley, 1953). This explanation method has shown to combine several desirable mathematical properties such as local accuracy, missingness and consistency. Furthermore this model-agnostic explanation approach allows us to use any MVPA decoding model in our workflow.
Increase in spatial specificity
We show that incorporating XAI to evaluate feature importance has the additional benefit of increasing spatial specificity in SL analysis. Solely analyzing prediction accuracy, as typically done in current studies applying SL decoding, only reveals if there is somewhere within the SL radius information present. Thus, maps of decoding accuracies do not exactly represent maps of informative voxels (Etzel et al., 2013). Incorporating SHAP as a post hoc test can help us to reconstruct the precise locations of voxels within a SL that are most important for decoding, what helps us to increase spatial specificity of SL analysis.
Experiments
@@
We first investigate general properties of the FuSL framework employing an artificial MRI dataset and study how different combinations of informative and uninformative data sources affect its decoding performance at different levels of noise, levels of spatial autocorrelations and levels signal inhomogeneities across samples. We further demonstrate that the integration of FC, ReHo and fALFF can enhance the prediction of pharmacological treatment of participants with rs-fMRI data from a pharmacological fMRI study (Wein et al., 2024). Thereby, the increase in decoding accuracy by integrating information across metrics allows us to identify additional regions that are significantly altered by treatment with the benzodiazepine alprazolam. Finally we show how SHAP can then be used to trace back the differential impact of alprazolam on the individual fMRI markers. Based on these applications we demonstrate how FuSL can provide a flexible and interpretable framework for data fusion in neuroimaging.
2 Materials and Methods
@@ 2.1.1 Artificial dataset
In the beginning we use an artificial dataset to investigate basic properties of FuSL when combining different data sources. On one side we want to study decoding performance of FuSL when signal or information is present in multiple imaging metrics or modalities simultaneously in the same location. On the other side we explore the behavior when we combine noisy non-informative data sources with informative ones in our FuSL. To study these scenarios we generate 3 artificial source signals, which could represent 3 different neuroimaging metrics or modalities. In this toy example the objective of the FuSL is, based on these 3 sources, to discriminate between two groups. We generated data for two groups, each containing 30 samples, by inducing artificial signals in one group, but not the other. Therefore we set signal values to 1 within a circular region of interest (ROI). In source 1 and 2 we simultaneously induce a signal in the left frontal cortex setting values within a 5 mm radius to 1 and -1 respectively (defined as ROI 1). As a third source we induced a signal in the left parietal cortex, setting values to 1 within a 5 mm radius (defined as ROI 2). We then generated 30 samples from these signals by sampling the signal amplitudes from a Gaussian normal distribution with standard deviations of $\sigma=1$ and $\sigma=2$, to obtain datasets with different amount of signal variability across samples, as typically observed in different clinical or experimental studies. We added random Gaussian noise to each sample with varying powers $P_{noise}=[1,10,20]$ resulting in datasets with different signal to noise ratios $S/N=[1,0.1,0.05]$ respectively. In this manner we created datasets for which signals in source 1 and 2 overlapped in ROI 1, and one signal was only expressed in ROI 2 in source 3. This allowed us to study in ROI 1 the impact of combining informative data sources, and in ROI 2 the impact of adding uninformative sources to informative ones. To further address the diversity in preprocessing protocols and modalities of data possibly integrated in FuSL, we additionally studied the impact of increasing the amount of spatial autocorrelations in artificial data. For this aim, we smoothed the data with a Gaussian filter with different full widths at half maximum $FWHM=[2mm,4mm,6mm]$ and evaluated decoding performance in dependency of the spatial smoothness. Finally, we also conducted a power analysis by sub-sampling from the whole artificial dataset and evaluated the proportion of vertices that exceeded statistical significance in dependence of the sample size.
@@ 2.1.2 Resting-state fMRI dataset
We used rs-fMRI data from 34 healthy male participants between the ages of 20 and 50 (mean = 27.2, std = 6.84) years and a BMI between 20.0 and 32.1 (mean = 24.3, std = 3.1). After screening for the absence of physical and psychiatric disorders all participants underwent rs-fMRI after a five-days intake of either alprazolam or placebo (Wein et al., 2024). Participants gave their informed written consent at the beginning of the experiment and the study was approved by the local ethics committee (approval number 18-1197-111) and the National Institute for Pharmaceutical Security (BfArM, Eudra-CT-number: 2018-002181-40). Subjects received alprazolam (1.5 mg/d in 3 doses of 0.5 mg) or placebo for 5 days, each with at least 7 days washout phases between treatments.
@@
For preprocessing the functional and structural imaging data, we incorporated the processing pipeline provided by fMRIPrep [^2] https://fmriprep.org/en/stable/ (version 20.2.4) (Esteban et al., 2017). Structural preprocessing included averaging T1w images across sessions, bias field correction based on Advanced Normalization Tool’s (ANTs) ‘N4BiasFieldCorrection’ (Tustison et al., 2010), brain tissue segmentation using FSL’s ‘fast’ (Jenkinson et al., 2012), brain surface reconstruction using FreeSurfer’s ‘recon-all’ (Fischl, 2012) and spatial normalization to the MNI152 standard space using ANT’s ‘antsRegistration’ (Avants et al., 2011). Functional preprocessing included susceptibility distortion correction based the double-echo gradient field maps using fMRIPrep’s costum workflows (Esteban et al., 2017). Functional images were registered to the anatomical volume using Freesurfer’s ‘bbregister’. Head motions were estimated usings FSL’s ‘flirt’ and slice time correction was applied using AFNI’s ‘3dTshift’ (Cox, 1996). Finally volumetric fMRI data were projected onto the cortical surface, generating high-resolution grayordinate time courses in the fsLR standard space (Glasser et al., 2013). Prior to the ReHo and fALFF analysis we smoothed the fMRI data using Gaussian surface smoothing with a FWHM of 3 mm (Glasser et al., 2013).
@@
Different rs-fMRI metrics allowed us to capture the different characteristics of the complex spatio-temporal dynamics of the BOLD signal. For analyzing local connectivity based on regional homogeneity (ReHo), we selected for each vertex its 5-hop neighboring vertices in the fsLR standard space (Glasser et al., 2013), and computed ReHo as the average Pearson correlation between neighbouring BOLD time courses within the 0.01 Hz - 0.1 Hz frequency range. Besides local connectivity, fractional amplitude of low frequency fluctuations (fALFF) can supplement such an analysis by detecting local modulations in the signals amplitude (Yang et al., 2007, Zou et al., 2008). This quantity reflects the ratio of the BOLD signal’s power spectrum of low frequencies to the signal’s entire frequency range. In our study we focused on the very-low frequency range 0.01 Hz - 0.05 Hz, which has been reported to increase during midazolam sedation and anesthesia (Kiviniemi et al., 2000, 2003, 2005). In addition to these local metrics, whole-brain functional connectivity (FC) allowed us to investigate modulations in brain networks from a more global perspective. Measures from graph theory thereby helped us to characterize changes in connectivity properties of individual regions of interest (ROIs). To evaluate whole-brain FC we subdivided each hemisphere into $180$ ROIs defined by Glasser et al. (2016). We computed the average activity time course within each ROI and filtered the resting-state BOLD signal within the 0.01 Hz - 0.1 Hz low frequency range. FC was then defined as the Pearson correlation coefficient $r$ between time courses of pairs of ROIs to obtain a $360\times 360$ whole-brain FC matrix. We selected a moderate FC threshold $r\geq 0.6$ yielding an average connection density of $39.2\%$ and analyzed modulations in functional connectivity efficiency (FCE) for each ROI in the network. According to the weights of the probabilistic multi-modal parcellation (Glasser et al., 2016), for each vertex we computed FCE as a weighted sum of the ROI-wise FCE values, weighting each vertex with the probability to belong to a respective ROI. This yielded a spatially continuous distribution of FCE values across vertices. A mathematical definition of these three metrics is additionally provided in supplement I. We computed each of these metrics for the alprazolam and placebo condition of each subject.
2.2 Fusion searchlight
In our FuSL framework, we integrated the complementary information provided by different data sources in a searchlight-based decoding analysis. For decoding pharmacological treatment of subjects we computed ReHo, fALFF and FCE at each vertex, and concatenated these values within each SL to train a classifier (figure 1 A and B). In this repeated measurement fMRI study we removed the within subject mean of each rs-fMRI metric to account for the treatment independent variability of metrics across subjects. In consideration of the variability in density of the fsLR standard space mesh (illustrated in supplementary figure S1), we defined SLs based on a k-hop neighborhood radius to ensure a constant number of features used to train each SL decoder across the cortex. We used cross validation (CV) with stratification to balance the classes within each split (Kohavi, 1995). To keep the test set completely independent we held out both fMRI sessions of each subject for testing at each CV split. After training and testing the SL classifier we computed a spatial map of test accuracies averaged across folds, pointing out vertices with neighborhoods that are informative to decode a specific brain state, like in classical SL analysis (figure 1 (C)). To identify test accuracies that were significantly different from chance level we permuted the labels 1000 times and applied threshold-free cluster enhancement (TFCE) to the original and permuted test accuracy maps (Smith and Nichols, 2009). To correct for multiple comparisons across vertices we took the maximum TFCE values of test accuracy maps derived from permuted data to build the null distribution and tested the values of the original TFCE map against them (Smith and Nichols, 2009). For the statistical testing of the decoding accuracy of different models across CV folds, we employed a paired t-test incorporating the variance correction proposed by Nadeau and Bengio (1999) to account for correlations in the training data across folds (Bouckaert and Frank, 2004). We applied false discovery rate (FDR) correction to account for multiple comparisons of different models (Benjamini and Hochberg, 1995). For the statistical analysis we utilized functions provided by the SciPy [^3] https://scipy.org/ and statsmodels [^4] https://www.statsmodels.org/stable/index.html Python packages (Virtanen et al., 2020, Seabold and Perktold, 2010).
In a next step we aimed to explain the predictions of the FuSL classifier. To reveal which metrics drove the decision of the SL, we used Shapley additive explanations (SHAP) to reconstruct the importance of each metric at each location on the cortex (Lundberg and Lee, 2017). This explanation approach assigns post-hoc an importance score to each input feature of a classifier, which we utilized to determine whether modulations in ReHo, fALFF or FCE were indicative for decoding a specific brain state. For each SL we first computed the feature importance of each metric at each vertex and then averaged these values across overlapping SLs at each vertex. In this manner we obtained one importance score for each metric at each vertex on the cortex. A formal derivation of these importance measure is provided in the supplementary material. To obtain maps that reflect simultaneously the importance and the informativeness of a metric, we further weighted spatial maps of SHAP values (feature importance) with the spatial map of decoding test accuracies (informativeness). We denoted this combination of informativeness and importance as the actual impact of a metric (figure 1 (D)). However, these maps do not provide us with information whether a metric increased or decreased due to an intervention or condition, which can be valuable information for the interpretation of analysis results. To address this shortcoming we propose to additionally incorporate a third characteristic of the data. We computed the differences in metrics between the groups across the cortex to obtain a spatial map indicating the intervention-associated changes of each metric. Finally we weighted these maps of feature differences with the feature impact. This yielded spatial maps that condensed the modulation, importance and informativeness of each metric at each location for a compact visual interpretation. We denote this combination as the feature-weighted impact. For a more specific analysis of these three aspects one can also visualize these maps individually, depending on the requirements of MRI study. An implementation including a demo version is available at: https://github.com/simonvino/FuSL which is based on functionalities provided by NiLearn [^5] https://nilearn.github.io/stable/index.html (Abraham et al., 2014), scikit-learn [^6] https://scikit-learn.org/stable/ (Pedregosa et al., 2011), and the SHAP software package [^7] https://shap.readthedocs.io/en/latest/ (Lundberg and Lee, 2017).